Which one of the following is true?(rounding to the nearest hundredths)


Question 22 

Which one of the following is true?(rounding to the nearest hundredths) 

  • You invest $100 today. If you earn 2% per quarter, the annual percentage rate is 6%. 

  • The effective annual rate of 6% compounded continuously is 6.4%. 

  • If the monthly rate is 0.2%, the annual percentage rate is 1.2% 

  • If the APR is 12% with daily compounding, effective annual rate is 12.75%

  • If the annual percentage rate is 12% with quarterly compounding, then the period rate is 4% per quarter.



Question 23 

Solve for the unknown number of years: Present Value-$810, Interest Rate=8%, Future Value=$1,821, Number of years=? 

  • 28.71 years 

  • 17.42 years 

  • 31.99 years 

  • 10.53 years 

  • 12.43 years
Answer #1

Q-22

The Answer is d)

As :-

=(1+12%/365)^365-1 = 12.75%

Q-23

Number of years have been calculated using the NPER function:-

=NPER(rate,pmt,pv,fv)

=NPER(8%,0,-810,1821)

=10.53 years